OpenStudy (anonymous):

Question from sis What is the least common multiple of 18 and 40? A. 2 B. 40 C. 180 D. 360

OpenStudy (anonymous):

i'm not allowed to answer for my sis

OpenStudy (anonymous):

To find the LCM, first factor the 2 numbers. Then find the product of all the factors they share and all the factors that are not shared.

OpenStudy (anonymous):

So 18 = 3*3*2 40 = 5*2*2*2

OpenStudy (anonymous):

So the only one they share is one of the 2's.

myininaya (myininaya):

or 18*1=18 14*1=14 18*2=36 14*2=28 18*3=54 14*3=42 18*4=74 14*4=56 18*5=90 14*5=70 18*6=108 14*6=84 18*7=126 14*7=98 18*8=146 14*8=112 18*9=162 14*9=126 so this can get pretty long but we see that the lcm(18,14)=126

OpenStudy (anonymous):

that's a complicated way to find it.

myininaya (myininaya):

lol and really long way sometimes

OpenStudy (anonymous):

The LCM is 5*3*3*2*2*2

myininaya (myininaya):

oh 40 darn

OpenStudy (anonymous):

The LCM of any set of numbers l and k is the product of A and B where A is the product of all the factors l and k have in common, and B is the product of all the factors l and k do not share.

myininaya (myininaya):

1*40=40 2*40=80 3*40=120 4*40=160 5*40=200 6*40=240 7*40=280 8*40=320 9*40=360 so yes the lcm(18,40)=360 doing it the long way lol

OpenStudy (anonymous):

thanks and that is way long

myininaya (myininaya):

you do really big numbers you might want to make a computer mathematical algorithm lol

OpenStudy (zarkon):

I would us the fact that LCM(x,y)=x*y/GCD(x,y)

myininaya (myininaya):

if you choose to do the long way that is polpak's way is much better

myininaya (myininaya):

thats nice zarkon

OpenStudy (zarkon):

use the euclidean algorithm to find the GCD

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Usually my way is faster than euclid's for the kinds of LCM's they have you look for in this level of math ;p

myininaya (myininaya):

40=18*2+4 18=4*4+2 4=2*2 ------------------------- gcd is 2

OpenStudy (zarkon):

sure...but for large numbers it could be hard to factor the number into primes

OpenStudy (anonymous):

Yep

myininaya (myininaya):

or looking at this we only see that they have 2 in common \[40=2*2*5, 18=2*3*3\]

myininaya (myininaya):

oops 40=2*2*2*5

myininaya (myininaya):

so using zarkon's formula we get lcm(18,40)=18*40/(gcd(18,40))=720/2=360

OpenStudy (anonymous):

Which was what I said to. ;p

OpenStudy (anonymous):

Err too

myininaya (myininaya):

thats right! polpak is always right of course lol

OpenStudy (anonymous):

Of course! ;p

myininaya (myininaya):

:)

OpenStudy (zarkon):

I just use my calculator, because I roll like that.

myininaya (myininaya):

thats a cool way to roll

OpenStudy (anonymous):

did you make a program for euclid's?

OpenStudy (zarkon):

I did...and to find the GCD as a linear combination

myininaya (myininaya):

can you write programs in your calculator?

myininaya (myininaya):

what kind do you have polpak?

OpenStudy (zarkon):

9*18+(-4)*40=2

OpenStudy (anonymous):

Ti 89

OpenStudy (zarkon):

ti-nspire cas

OpenStudy (anonymous):

nice

myininaya (myininaya):

i do see word prgm on my calculator maybe you can write programs in ti83

OpenStudy (zarkon):

I wrote the program a few weeks ago then I find out that it has the program built in ...oh well ;)

myininaya (myininaya):

lol writing programs is fun!

OpenStudy (zarkon):

Most of my programming is done in Matlab though

myininaya (myininaya):

i like maple

OpenStudy (zarkon):

to program in...never was a big fan.

OpenStudy (zarkon):

I love what it can do, but I don't like the programming language

myininaya (myininaya):

i'm guess i'm just use to it i still don't know everything about it i try to implement the rsa with it

OpenStudy (anonymous):

I like python

myininaya (myininaya):

i heard of matlab but not python i might have to look into that

OpenStudy (zarkon):

I'm just started programming in python...seems to be nice

OpenStudy (anonymous):

It is quite nice. Certainly my favorite language

OpenStudy (anonymous):

you guys are awesome o.O

OpenStudy (zarkon):

We know that ;)

myininaya (myininaya):

lol joe is incredible